AnimeGhoul8863:

Harrison and Sherrie are making decisions about their bank accounts. Harrison wants to deposit $200 as a principle amount, with an interest of 2% compounded quarterly. Sherrie wants to deposit $200 as the principle amount, with an interest of 4% compounded monthly. Explain which method results in more money after 2 years. Show all work.

BenLindquist:

\[\sqrt{2}\]

umm:

@BenLindquist - Where is your work?

AnimeGhoul8863:

To determine which method earns more money after 2 years we have to lay out our data. P=Principle which in this case equals $200 R=The Interest rate in this case the interest rate is 2% or 0.02 N= Will represent the number of times Sherrie or Harrison compounded the money per year in this case it equals 4 T= Is simply going to represent the Time it takes to earn the money in this case that is 2 years We now check each person to see how much they have earned. We are going to be using the compound Interest Formula: A=P(1+r/n)^nt Harrison: When Harrison first starts he is going to make (1+0.02/4)^4 = 1.02015 but after 2 years go by his money is going to increase and he is going to earn ($200)(1.02015)^2 = $208.14 after 2 years. Sherrie: When Sherrie first starts she is going to make (1+0.04/12)^12 = 1.04074 this is already more than what Harrison made when he first started.But after 2 years Sherrie's money is going to increase and she is going to earn ($200)(1.04074)^2 = $216.63 this means Sherrie has earned 8 more dollars than Harrison making her method correct and effective.

AnimeGhoul8863:

my answer lmao Benneh no that isnt right e.e